OpenStudy (anonymous):

The perimeter of a triangle with 2 equal sides is 42 cm. If its base were lengthened by 7 cm and each leg were shortened by 5 cm, all 3 sides would be equal. Find the length of the base the original triangle.

OpenStudy (dumbcow):

base = 6

OpenStudy (dumbcow):

2x+b = 42 b+7 = x-5 solve for b

OpenStudy (anonymous):

You need to solve simultaneous equations. For the 1st triangle, let the two equal sides be x, and the base be y, then you get 2x+y = 42 for the 2nd triangle, the base is y+7, and the two equal sides are each x-5. Then you're also told that the 2nd one is equilateral, that is y+7 = x-5 So equation 1: 2x+y=42 Equation 2: x-y=12 ( after rearranging the y+7=x-5 )

OpenStudy (lgbasallote):

let x = leg y = base 42 = 2x + y 3x = y+7 + 2(x - 5) 3x = y + 7 + 2x - 10 x = y - 3 2x + y = 42 x - y = -3 3x = 39 x = 13 y = 16 cm is the orig base